# Momentum dependent quantum Ruelle-Pollicott resonances in translationally invariant many-body systems

### v2

@@ Abstract
We study Ruelle-Pollicott resonances in translationally invariant quantum many-body lattice systems via spectra of momentum-resolved operator propagator on infinite systems. Momentum dependence gives insight into decay of correlation functions, showing that, depending on their symmetries, different correlation functions in general decay with different rates. Focusing on the kicked Ising model the spectrum seems to be typically composed of an annular random matrix like ring whose size we theoretically predict, and few isolated resonances. We identify several interesting regimes, including a mixing regime with a power-law decay of correlation functions. In that regime we also observe a huge difference in time-scales of different correlation functions due to an almost conserved operator. An exact expression for the singular values of the operator propagator is conjectured, showing that it becomes singular at a special point.
@@ I Introduction
As is often the case tough, theoretical elegance comes with a price. It is not clear how to in practice do the required “analytical continuation”. Mathematics is heavy, and even in exactly solvable classical cases like the cat map antoniou97 , the Bernoulli map or the baker map Hasegawa92 ; Tasaki93 (see Ref. gaspard95 for differential equations), one often rather starts by writing the spectral decomposition of $U$ on an appropriate system-dependent functional space. Thereby, the crux of the problem shifts to the question of what is the appropriate functional space.
@@
A more practical approach, and the only viable option in numerical calculations of RP resonances, is to make a non-unitary operator $U(\epsilon)$ out of a unitary $U$ by adding some “dissipation” and then identifying RP resonance as isolated eigenvalues in the limit of zero dissipation strength $\epsilon$. Dissipation is often added in the form of a noise or doing some coarse graining. Such a noisy or a coarse graining procedure adds a practical appeal to an otherwise rather mathematical formulation of RP resonances. In real life we are never really in the unitary limit – the system of interest is always invariably coupled to some extra degrees of freedom, and/or, in a many-body (quantum) system we never have access to all observables and therefore full unitary evolution. If we study system numerically we also necessarily have numerical errors, meaning that we are again studying a noisy non-unitary system. From a practical point of view an exact unitarity is therefore just a mirage. Non-unitary prescription to get the RP resonances is exactly what one anyway has to deal with. Let us finish this introduction with a few words about the spectral decomposition of such non-unitary linear operators.
@@
As we have mentioned, for some classical solvable maps Hasegawa92 ; Tasaki93 ; antoniou97 one can explicitly calculate the spectral decomposition. From those solutions one sees that the eigenvectors are rather singular objects – they involve Dirac delta functions and their derivatives – and are therefore “too wild” to be in $L^{2}$. They are not really smooth densities but rather functionals. The space on which these functionals are allowed to act also excludes for instance Dirac delta functions, meaning that the intricacies of mathematical formalism actually forces us to avoid asking meaningless questions, like about individual trajectories in a chaotic unstable system, and instead focus on evolution of smooth densities (in quantum mechanics on evolution of density operators rather than of pure states). One can also reach the same conclusion about the singular behavior of spectral decomposition in a more intuitive way: a common scenario of chaos in smooth classical systems is the so-called stretch-and-fold mechanics, see e.g. Ref. chaosbook ; ott , due to stable and unstable manifolds. As one evolves an initial smooth density $\rho(0)$ it is stretched along the unstable manifold. Therefore, $\rho(t)$ becomes smoother along the unstable manifold, and at the same time less smooth along the stable manifold. At long times $\rho(t)$ will therefore become localized around the unstable manifold. With the spectral decomposition in mind this suggests that the right eigenvectors ${|R_{j}\rangle}$ will be localized along the unstable manifold, while the left ones ${|L_{j}\rangle}$ will be localized along the stable manifold. Because in a chaotic system those manifolds have a fractal structure with details on an arbitrarily small scale, one expect similar singular (fractal) behavior also in eigenvectors. That is fine though, again, in accordance with chaotic systems having positive rate of generating information (as measured e.g. by dynamical entropies ott ; gaspard ), this high complexity is apparently hidden in eigenvectors, which, though, on themselves do not have much “meaning” – only their action on smooth initial densities is what is physically observable. Physical information about relaxation rates toward equilibrium is on the other hand simply encoded in eigenvalues $\lambda_{j}$ foot5 . This splitting between information contained in eigenvectors and eigenvalues is less clear if one instead looks on a unitary propagator $U$.
@@ I.1 Previous work
Exact results on RP resonances are possible for classical hyperbolic maps, like the baker and the Bernoulli map Hasegawa92 ; Tasaki93 , and the cat map antoniou97 . The standard map has been studied in Refs. fishman00 ; venegeroles08 ; mata23 , while the kicked top, including a regime where classical dynamics is mixed, in Ref. weber . A detailed analysis of the multibaker map is in Ref. gaspard92 ; gaspard , while coupled baker maps have been studied in Refs. arul94 ; Haake03 . Rather than working with a propagator one can approach everything from the point of view of spectral analysis of correlation functions, extending frequencies to the complex domain gaspard . Thereby one extracts RP resonances directly from correlations, see Ref. martin for an example. More rigorous treatment and a comparison between classical and quantum cases with noise for maps on a torus can be found in Ref. Nonnenmacher . Another technical trick of how to obtain RP resonances is using periodic orbits with trace formulas and zeta functions chaosbook ; sridhar ; altshuler ; leboeuf . It has been observed that sometimes there might be issues with convergence of RP resonances for particular choices of coarse-graining blank98 ; shepelyansky ; shudo21 (could perhaps be related to strong non-normality of matrices).
In single-particle quantum chaos Haake spectral criteria of quantum chaos, e.g., nearest-neighbor level spacing, proved to be useful. In many-body systems the mean level spacing is exponentially small and therefore in some situations spectral quantum chaos criteria at smallest energy scales may be inappropriate. The RP formalism instead is a setting that can be used. Quantized perturbed cat map with noisy shift being described by a Kraus map has been studied in Ref. mata0304 . More recently it has been proposed Mori to use Lindblad dissipation, calculating the Lindbadian gap in an appropriate limit $\lim_{\epsilon\to 0}\lim_{L\to\infty}$ (note that this is the difficult order of limits to evaluate as one usually has to get the gap in the thermodynamic limit at a finite $\epsilon$), see also results in Ref. sa that can be used to that end. Last two works study many-body quantum systems, in particular Ref. Mori the kicked Ising (KI) model whose RP resonances have been studied for the first time already more than two decades ago in Refs Prosen ; Prosen07 .
@@ I.2 New results
We approach RP resonances by studying the spectrum of a propagator of operators on an infinite system. Resolving the (quasi)momentum dependence and truncating it down to operators with support on $r$ consecutive sites, we can discuss decay of any correlation function. To illustrate our results we are going to study the same kicked Ising model and use essentially the same truncation method used in Ref. Prosen , but extend it from the zero momentum subspace studied there. Doing that we will find many new and interesting results. As we work directly in the thermodynamic limit, i.e., study evolution of operators in an infinite system, there is only one limit we need to take, namely, the operator support $r$ to infinity, instead of typically two for other types of coarse-graining or noise approaches to RP resonances. We show that to properly assess dynamical properties of the system it pays off to take into account the momentum dependence of RP resonances because different observables in general decay with different rates. The spectrum of eigenvalues in the KI model consists of a random-matrix like bulk of an annular shape whose radii we analytically predict, and few isolated resonances. Those indicate rich physics: besides identifying expected and known chaotic regime, we also find a mixing non-chaotic regime with a power-law decay of correlations due to a branch cut. This regime is rather interesting because, despite being away from any integrable point, different observables decay on vastly different timescales. For instance, for a particular choice of parameters magnetization correlations start to decay on a $\sim 10^{4}$ times longer time scale than for instance the staggered magnetization or magnetization current correlations. The origin of such disparate timescales is due to an almost conserved operators giving rise to a prethermalization plateau. Using analytical expression for singular values of the truncated propagator we also conjecture a lower bound on RP resonances that depends only on the parameter $\tau$. Singular values also highlight speciality of KI parameters when $\tau$ is an odd multiple of $\pi/4$, for which some singular values become 00 (this point includes an even more special dual unitary KI model DU if one in addition sets $h_{\rm x}=1$).
@@ I.3 Kicked Ising model
There are three parameters $\tau$, $h_{\rm x}$, and $h_{\rm z}$ (we use the same parameterization as in Ref. Mori , see Appendix A for an alternative used in Ref. Prosen ). For generic parameters the model is nonintegrable without any local conserved quantities. We measure time in units of $\tau$, i.e., our $t$ takes integer values.
The KI model is integrable at special values of parameters. Known integrable points are (i) $\tau h_{\rm x}=0$ (or an integer multiple of $\pi$) for which the Hamiltonian is trivially diagonal in the computational basis, (ii) $\tau h_{\rm z}=0$ (or an integer multiple of $\pi$), i.e., transversal field, for which it is integrable prosenPTPS , similar to integrability of the autonomous transversal Ising model, (iii) integrable points when the strength of the field $\gamma$ defined in Eq.(28) is a multiple of $\pi/2$ Prosen07 . This is achieved when either $\tau h_{\rm x}$ or $\tau h_{\rm z}$ is an odd multiple of $\pi/2$, or when both $\tau h_{\rm x}$ and $\tau h_{\rm z}$ are integer multiples of $\pi$ (for our choice of $h_{\rm x}=0.9$ and $h_{\rm z}=0.8$ two integrable points with the smallest $\tau$ are $\tau\approx 1.75$ and $\tau\approx 1.96$, shown with black points in Fig. 2). New trivial integrable point (iv) is when $\tau=\pi/2$ (or an odd multiple of $\pi/2$; the leftmost black point in Fig. 2) and arbitrary $h_{\rm x}$ and $h_{\rm z}$, for which the model is noninteracting because ${\rm e}^{{\rm i}\sigma^{\rm z}_{1}\sigma^{\rm z}_{2}\pi/2}={\rm i}\sigma^{\rm z}_{1}\sigma^{\rm z}_{2}$, and therefore the whole interacting part is trivial $\prod_{j}{\rm e}^{{\rm i}\tau\sigma^{\rm z}_{j}\sigma^{\rm z}_{j+1}}=({\rm i})^{L}\mathbbm{1}$ (for PBC). Some other non-integrable special points will be mentioned in Sec. II.2. The propagator $U$ has several appealing properties (which was one of the reasons it was chosen in Ref. Prosen ): it can be viewed as a quantum circuit made of commuting 1-qubit $\sigma^{\rm z}$ gates, then commuting 2-qubit $\sigma^{\rm z}_{j}\sigma^{\rm z}_{j+1}$ gates, followed finally by again commuting 1-qubit $\sigma^{\rm x}$ gates. The fact that it has a single layer of commuting 2-qubit gates means that the operator support can increase by at most $2$ sites in a single step. For instance, propagating a translationally invariant (TI) operator with support on $r=3$ sites will after one step result in operators with support on at most $r^{\prime}=5$ sites, e.g.,
@@
The main object we will study will be autocorrelation functions $C(t)$ of different extensive observables $A=\sum_{j}a_{j}{\rm e}^{{\rm i}kj}$,
@@
Autocorrelation function $C(t)$ will be calculated in three different ways: (i) evaluating Eq.(3) in a finite system with periodic boundary conditions ($L\leq 33$) by exactly evolving random initial state, (ii) approximating it by a truncated operator propagator on an infinite system, and (iii) by calculating the leading RP resonance of the truncated propagator, giving the asymptotic decay $\asymp|\lambda_{1}|^{t}$.
@@ II.1 Truncated propagator of translated local operators
in short, the Pauli basis. The operator basis on $r$ sites is then formed by a direct products of operators from the Pauli basis on different sites $j$. Let us classify local operators according to their maximal support size $r$. The set of operators ${\cal S}_{r}$ is made of all operators that start with a non-identity operator on site $1$, i.e., one in the set $\mathcal{p}_{1}=\{\sigma^{\rm x}_{1},\sigma^{\rm y}_{1},\sigma^{\rm z}_{1}\}$, and have the last non-identity operator at most at the site $r$. More precisely, ${\cal S}_{r}$ is a set of all operators
@@
The translation operator by one site is denoted by $T$, and acts like e.g. $T(\sigma^{\rm x}_{1}\sigma^{\rm z}_{2})=\sigma^{\rm x}_{2}\sigma^{\rm z}_{3}$. The operator basis can be organized according to the momentum $k$ that labels eigenvalues ${\rm e}^{{\rm i}k}$ of $T$. In a finite system $k$ takes a discrete set of values, while in the infinite lattice it is a continuous variable $k\in[0,2\pi)$. The $k$-resolved basis with support on at most $r$ sites is therefore composed of operators $A$ that are a translational sum of a local operator $a$,
@@
Because $H$ commutes with $T$, it has a block structure and one can treat dynamics in each $k$-subspace separately. Unitary Floquet one-step propagator $U$ induces evolution on the space of operators as $O^{\prime}=U^{\dagger}OU$. Such evolution is unitary on the whole infinite dimensional operator space. However, if one truncates the operator space to translational sums of operators from ${\cal S}_{r}$, like those in Eq. (8), one ends up with a nonunitary truncated propagator $M(k)$. Eigenvalues of $M(k)$ will be our RP resonances that we are interested in. Note that $M(k)$ directly acts on an infinite system and so the only remaining parameter is the support size $r$ that also determines the size $N$ (7) of the matrix $M(k)$.
For our KI system (I.3) constructing $M(k)$ is easy. As we explained, one application of $U$ can increase the support of $A$ only by one site at each edge of $a$. Therefore, to get the matrix element $[M(k)]_{A,A^{\prime}}$ between two translational operators $A$ and $A^{\prime}$, defined in terms of local $a$ and $a^{\prime}$ from ${\cal S}_{r}$ (8), respectively, it suffices to act with $U$ on an operator supported on $r+2$ sites. That is, starting with $a^{\prime}\in{\cal S}_{r}$ supported on sites $1,\ldots,r$ we extend it to $r+2$ sites by concatenating it with two identities on each end to accommodate for a possible spreading of $a^{\prime}$ at each end by at most one site, and then transform it with $U$. The resulting transformed $a^{\prime}$ can be expanded in terms of operators $b$ on $r+2$ sites,
@@
Compared to other possible ways of obtaining RP resonances, like using noise or some coarse-graining, this truncation procedure has several appealing features Prosen . It is physically motivated as one is usually interested in local operators, and, because one directly works in the thermodynamic limit, one has to take only one limit, that is $r\to\infty$, instead of usually two like the noise strength $\epsilon\to 0$ as well as the system size $L\to\infty$. The matrix $M(k)$ is also equal to a truncation of a unitary propagator in the $k$ sector of a finite system with PBC, provided $L\geq 2r-1$ (i.e., taking a finite system with PBC one selects in a given momentum block only operators with support on $\leq r$ sites). Therefore one can consider our infinite system size $M(k)$ as a specific operator support-selective way of truncating an appropriate unitary matrix. Let us mention at this point that within a random matrix theory one has studied truncations of Haar random unitary matrices, and they result in a spectrum inside a circle whose radius depends on the fraction of retained vectors karol00 .
When numerically dealing with $M(k)$ we never really write out the full matrix (its size $N$ gets too large for large $r$, e.g., $N\approx 5\cdot 10^{7}$ for $r=13$), but rather generate it on the fly by evaluating Eq.(9). This is done by applying each gate of the quantum circuit representation of $U$, see Appendix B.
Due to translational symmetry one also has that $M(-k)$ has the same spectrum as $M(k)$ and therefore one can limit oneself to $k\in[0,\pi)$, rather than $k\in[0,2\pi)$ (spectrum at $2\pi-k$ is the same as at $k$). In addition to that, due to a spatial reflection symmetry (reversing the order of operators, i.e., in a finite system corresponding to sites mapping $j\to L-j$), which the KI Hamiltonian (I.3) has, the $k=0$ sector splits into two independent parity subsectors corresponding to eigenvalues ±$\pm 1$ of the reflection operator. Namely, $N$ elements of the local basis ${\cal S}_{r}$ can be split into two sets according to whether the reflected operator $Ra$ is the same as $a$, where the reflection of the local $a$ is defined by
@@
with $p$ being the index of the rightmost operator that is not equal to identity, i.e., $a_{p}\in\mathcal{p}$ but $a_{p+1,\ldots,r}=\mathbbm{1}$. That is ${\cal S}_{r}={\cal S}_{r}^{\rm pairs}\cup{\cal S}_{r}^{\rm single}$, where ${\cal S}_{r}^{\rm single}=\{a|(a\in{\cal S}_{r})\,\land\,(Ra=a)\}$, and ${\cal S}_{r}^{\rm pairs}=\{a|(a\in{\cal S}_{r})\,\land\,(Ra\neq a)\}$. Instead of using the basis given by the elements of ${\cal S}_{r}$, one takes an odd basis $\{(a-Ra)/\sqrt{2}|a\in{\cal S}_{r}^{\rm pairs}\}$, and an even basis $\{a|a\in{\cal S}_{r}^{\rm single}\}\cup\{(a+Ra)/\sqrt{2}|a\in{\cal S}_{r}^{\rm pairs}\}$ (where each pair is of course accounted for only once and not twice). The corresponding even and odd sector matrices $M(k=0^{+})$ and $M(k=0^{-})$ are then constructed as in Eqs.(9) and (10) with the only difference being that $a^{\prime}$ and $a$ run over elements of the even or the odd basis, instead of over full ${\cal S}_{r}$.
As we shall see, studying the momentum dependent RP resonances will give us much more information that just limiting oneself to $k=0^{+}$ (a sector to which e.g. total magnetization belongs). In particular, different dynamical regimes, like chaotic, mixing, and integrable, will be easier to identify. Note also that some important frequently studied operators are not from the $k=0^{+}$ sector, an example being the magnetization current (i.e., charge current in fermionic language) operator $J$ (the KI model of course does not conserve the magnetization)

### v3

@@ Abstract
We study Ruelle-Pollicott resonances in translationally invariant quantum many-body lattice systems via spectra of a momentum-resolved operator propagator on infinite systems. Momentum dependence gives insight into the decay of correlation functions, showing that, depending on their symmetries, different correlation functions in general decay with different rates. Focusing on the kicked Ising model the spectrum seems to be typically composed of an annular random matrix like ring whose size we theoretically predict, and few isolated resonances. We identify several interesting regimes, including a mixing regime with a power-law decay of correlation functions. In that regime we also observe a huge difference in time-scales of different correlation functions due to an almost conserved operator. An exact expression for the singular values of the operator propagator is conjectured, showing that it becomes singular at a special point.
@@ I Introduction
As is often the case tough, theoretical elegance comes at a price. It is not clear how to in practice do the required “analytical continuation”. Mathematics is heavy, and even in exactly solvable classical cases like the cat map antoniou97 , the Bernoulli map or the baker map Hasegawa92 ; Tasaki93 (see Ref. gaspard95 for differential equations), one often rather starts by writing the spectral decomposition of $U$ on an appropriate system-dependent functional space. Thereby, the crux of the problem shifts to the question of what is the appropriate functional space.
@@
A more practical approach, and the only viable option in numerical calculations of RP resonances, is to make a non-unitary operator $U(\epsilon)$ out of a unitary $U$ by adding some “dissipation” and then identifying RP resonances as isolated eigenvalues in the limit of zero dissipation strength $\epsilon$. Dissipation is often added in the form of a noise or doing some coarse graining. Such a noisy or a coarse graining procedure adds a practical appeal to an otherwise rather mathematical formulation of RP resonances. In real life we are never really in the unitary limit – the system of interest is always invariably coupled to some extra degrees of freedom, and/or, in a many-body (quantum) system we never have access to all observables and therefore full unitary evolution. If we study a system numerically we also necessarily have numerical errors, meaning that we are again studying a noisy non-unitary system. From a practical point of view an exact unitarity is therefore just a mirage. Non-unitary prescription to get the RP resonances is exactly what one anyway has to deal with. Let us finish this introduction with a few words about the spectral decomposition of such non-unitary linear operators.
@@
As we have mentioned, for some classical solvable maps Hasegawa92 ; Tasaki93 ; antoniou97 one can explicitly calculate the spectral decomposition. From those solutions one sees that the eigenvectors are rather singular objects – they involve Dirac delta functions and their derivatives – and are therefore “too wild” to be in $L^{2}$. They are not really smooth densities but rather functionals. The space on which these functionals are allowed to act also excludes for instance Dirac delta functions, meaning that the intricacies of mathematical formalism actually forces us to avoid asking meaningless questions, such as about individual trajectories in a chaotic unstable system, and instead focus on the evolution of smooth densities (in quantum mechanics on the evolution of density operators rather than of pure states). One can also reach the same conclusion about the singular behavior of spectral decomposition in a more intuitive way: a common scenario of chaos in smooth classical systems is the so-called stretch-and-fold mechanism, see e.g. Refs. chaosbook ; ott , due to stable and unstable manifolds. As one evolves an initial smooth density $\rho(0)$ it is stretched along the unstable manifold. Therefore, $\rho(t)$ becomes smoother along the unstable manifold, and at the same time less smooth along the stable manifold. At long times $\rho(t)$ will therefore become localized around the unstable manifold. With the spectral decomposition in mind this suggests that the right eigenvectors ${|R_{j}\rangle}$ will be localized along the unstable manifold, while the left ones ${|L_{j}\rangle}$ will be localized along the stable manifold. Because in a chaotic system those manifolds have a fractal structure with details on an arbitrarily small scale, one expect similar singular (fractal) behavior also in eigenvectors. That is fine though, again, in accordance with chaotic systems having a positive rate of generating information (as measured e.g. by dynamical entropies ott ; gaspard ), this high complexity is apparently hidden in eigenvectors, which, though, on themselves do not have much “meaning” – only their action on smooth initial densities is what is physically observable. Physical information about relaxation rates toward equilibrium is on the other hand simply encoded in eigenvalues $\lambda_{j}$ foot5 . This splitting between information contained in eigenvectors and eigenvalues is less clear if one instead looks on a unitary propagator $U$.
@@ I.1 Previous work
Exact results on RP resonances are possible for classical hyperbolic maps, like the baker and the Bernoulli map Hasegawa92 ; Tasaki93 , and the cat map antoniou97 . The standard map has been studied in Refs. fishman00 ; venegeroles08 ; mata23 , while the kicked top, including a regime where classical dynamics is mixed, in Ref. weber . A detailed analysis of the multibaker map is in Ref. gaspard92 ; gaspard , while coupled baker maps have been studied in Refs. arul94 ; Haake03 . Rather than working with a propagator one can approach everything from the point of view of spectral analysis of correlation functions, extending frequencies to the complex domain gaspard . Thereby one extracts RP resonances directly from correlations, see Ref. martin for an example. More rigorous treatment and a comparison between classical and quantum cases with noise for maps on a torus can be found in Ref. Nonnenmacher . Another technical trick of how to obtain RP resonances is using periodic orbits with trace formulas and zeta functions chaosbook ; sridhar ; altshuler ; leboeuf . It has been observed that sometimes there might be issues with convergence of RP resonances for particular choices of coarse-graining blank98 ; shepelyansky ; shudo21 (which could perhaps be related to strong non-normality of matrices).
In single-particle quantum chaos Haake spectral criteria of quantum chaos, e.g., nearest-neighbor level spacing, proved to be useful. In many-body systems the mean level spacing is exponentially small and therefore in some situations spectral quantum chaos criteria at the smallest energy scales may be inappropriate. The RP formalism instead is a setting that can be used. The quantized perturbed cat map with noisy shift being described by a Kraus map has been studied in Ref. mata0304 . More recently it has been proposed Mori to use Lindblad dissipation, calculating the Lindbadian gap in an appropriate limit $\lim_{\epsilon\to 0}\lim_{L\to\infty}$ (note that this is the difficult order of limits to evaluate as one usually has to get the gap in the thermodynamic limit at a finite $\epsilon$), see also results in Ref. sa that can be used to that end. Last two works study many-body quantum systems, in particular Ref. Mori the kicked Ising (KI) model whose RP resonances were studied for the first time already more than two decades ago in Refs Prosen ; Prosen07 .
@@ I.2 New results
We approach RP resonances by studying the spectrum of a propagator of operators on an infinite system. Resolving the (quasi)momentum dependence and truncating it down to operators with support on $r$ consecutive sites, we can discuss the decay of any correlation function. To illustrate our results we are going to study the same kicked Ising model and use essentially the same truncation method used in Ref. Prosen , but extend it from the zero momentum subspace studied there. Doing that we will find many new and interesting results. As we work directly in the thermodynamic limit, i.e., study the evolution of operators in an infinite system, there is only one limit we need to take, namely, the operator support $r$ to infinity, instead of typically two for other types of coarse-graining or noise approaches to RP resonances. We show that to properly assess the dynamical properties of the system it pays to take into account the momentum dependence of RP resonances because different observables in general decay with different rates. The spectrum of eigenvalues in the KI model consists of a random-matrix like bulk of an annular shape whose radii we analytically predict, and few isolated resonances. Those indicate rich physics: besides identifying expected and known chaotic regime, we also find a mixing non-chaotic regime with a power-law decay of correlations due to a branch cut. This regime is rather interesting because, despite being away from any integrable point, different observables decay on vastly different timescales. For instance, for a particular choice of parameters magnetization correlations start to decay on a $\sim 10^{4}$ times longer time scale than for instance the staggered magnetization or magnetization current correlations. The origin of such disparate timescales is due to an almost conserved operator giving rise to a prethermalization plateau. Using an analytical expression for singular values of the truncated propagator we also conjecture a lower bound on RP resonances that depends only on the parameter $\tau$. Singular values also highlight the speciality of KI parameters when $\tau$ is an odd multiple of $\pi/4$, for which some singular values become 00 (this point includes an even more special dual unitary KI model DU if one in addition sets $h_{\rm x}=1$).
@@ I.3 Kicked Ising model
There are three parameters $\tau$, $h_{\rm x}$, and $h_{\rm z}$ (we use the same parameterization as in Ref. Mori , see Appendix A for an alternative used in Ref. Prosen ). For generic parameters the model is nonintegrable without any local conserved quantities. We measure time in units of $\tau$, i.e., our $t$ takes integer values with the propagator being $U^{t}$.
The KI model is integrable at special values of parameters. Known integrable points are (i) $\tau h_{\rm x}=0$ (or an integer multiple of $\pi$) for which the Hamiltonian is trivially diagonal in the computational basis, (ii) $\tau h_{\rm z}=0$ (or an integer multiple of $\pi$), i.e., transversal field, for which it is integrable prosenPTPS , similar to integrability of the autonomous transversal Ising model, (iii) integrable points when the strength of the field $\gamma$ defined in Eq.(28) is a multiple of $\pi/2$ Prosen07 . This is achieved when either $\tau h_{\rm x}$ or $\tau h_{\rm z}$ is an odd multiple of $\pi/2$, or when both $\tau h_{\rm x}$ and $\tau h_{\rm z}$ are integer multiples of $\pi$ (for our choice of $h_{\rm x}=0.9$ and $h_{\rm z}=0.8$ two integrable points with the smallest $\tau$ are $\tau\approx 1.75$ and $\tau\approx 1.96$, shown with black points in Fig. 2). New trivial integrable point (iv) is when $\tau=\pi/2$ (or an odd multiple of $\pi/2$; the leftmost black point in Fig. 2) and arbitrary $h_{\rm x}$ and $h_{\rm z}$, for which the model is noninteracting because ${\rm e}^{{\rm i}\sigma^{\rm z}_{1}\sigma^{\rm z}_{2}\pi/2}={\rm i}\sigma^{\rm z}_{1}\sigma^{\rm z}_{2}$, and therefore the whole interacting part is trivial $\prod_{j}{\rm e}^{{\rm i}\tau\sigma^{\rm z}_{j}\sigma^{\rm z}_{j+1}}=({\rm i})^{L}\mathbbm{1}$ (for PBC). Some other non-integrale special points will be mentioned in Sec. II.2.
The propagator $U$ has several appealing properties (which was one of the reasons it was chosen in Ref. Prosen ): it can be viewed as a quantum circuit made of commuting 1-qubit $\sigma^{\rm z}$ gates, then commuting 2-qubit $\sigma^{\rm z}_{j}\sigma^{\rm z}_{j+1}$ gates, followed finally by again commuting 1-qubit $\sigma^{\rm x}$ gates. The fact that it has a single layer of commuting 2-qubit gates means that the operator support can increase by at most $2$ sites in a single step. For instance, propagating a translationally invariant (TI) operator with support on $r=3$ sites will after one step result in operators with support on at most $r^{\prime}=5$ sites, e.g.,
@@
The main physical quantity that we study will be autocorrelation functions $C(t)$ of different extensive observables $A=\sum_{j}a_{j}{\rm e}^{{\rm i}kj}$,
@@
Autocorrelation function $C(t)$ will be calculated in three different ways: (i) evaluating Eq.(3) in a finite system with periodic boundary conditions ($L\leq 33$) by an exactly evolving random initial state, (ii) approximating it by a truncated operator propagator on an infinite system, and (iii) calculating the leading RP resonance of the truncated propagator, giving the asymptotic decay $\asymp|\lambda_{1}|^{t}$.
@@ II.1 Truncated propagator of translated local operators
in short, the Pauli basis. The operator basis on $r$ sites is then formed by a direct product of operators from the Pauli basis on different sites $j$. Let us classify local operators according to their maximal support size $r$. The set of operators ${\cal S}_{r}$ is made of all operators that start with a non-identity operator on site $1$, i.e., one in the set $\mathcal{p}_{1}=\{\sigma^{\rm x}_{1},\sigma^{\rm y}_{1},\sigma^{\rm z}_{1}\}$, and have the last non-identity operator at most at the site $r$. More precisely, ${\cal S}_{r}$ is a set of all operators
@@
The translation (super)operator by one site is denoted by $T$, and acts like e.g. $T(\sigma^{\rm x}_{1}\sigma^{\rm z}_{2})=\sigma^{\rm x}_{2}\sigma^{\rm z}_{3}$. The operator basis can be organized according to the momentum $k$ that labels eigenvalues ${\rm e}^{{\rm i}k}$ of $T$. In a finite system $k$ takes a discrete set of values, while in the infinite lattice it is a continuous variable $k\in[0,2\pi)$. The $k$-resolved basis with support on at most $r$ sites is therefore composed of operators $A$ that are a translational sum of a local operator $a$,
@@
Because $H$ commutes with $T$, it has a block structure and one can treat dynamics in each $k$-subspace separately. Unitary Floquet one-step propagator $U$ induces evolution on the space of operators as $O^{\prime}=U^{\dagger}OU$. Such evolution is unitary on the whole infinite dimensional operator space. However, if one truncates the operator space to translational sums of operators from ${\cal S}_{r}$, like those in Eq. (8), one ends up with a nonunitary truncated propagator $M(k)$. Eigenvalues of $M(k)$ will be our RP resonances in which we are interested. Note that $M(k)$ directly acts on an infinite system and so the only remaining parameter is the support size $r$ that also determines the size $N$ (7) of the matrix $M(k)$.
For our KI system (I.3) constructing $M(k)$ is easy. As we explained, one application of $U$ can increase the support of $A$ only by one site at each edge of $a$. Therefore, to get the matrix element $[M(k)]_{A,A^{\prime}}$ between two translational operators $A$ and $A^{\prime}$, defined in terms of local $a$ and $a^{\prime}$ from ${\cal S}_{r}$ (8), respectively, it suffices to act with $U$ on an operator supported on $r+2$ sites. That is, starting with $a^{\prime}\in{\cal S}_{r}$ supported on sites $1,\ldots,r$ we extend it to $r+2$ sites by concatenating it with identities on each end to accommodate for a possible spreading of $a^{\prime}$ at each end by at most one site, and then transform it with $U$. The resulting transformed $a^{\prime}$ can be expanded in terms of operators $b$ on $r+2$ sites,
@@
Compared to other possible ways of obtaining RP resonances, like using noise or some coarse-graining, this truncation procedure has several appealing features Prosen . It is physically motivated as one is usually interested in local operators, and, because one directly works in the thermodynamic limit, one has to take only one limit, that is $r\to\infty$, instead of usually two like the noise strength $\epsilon\to 0$ as well as the system size $L\to\infty$. The matrix $M(k)$ is actually equal to a truncation of a unitary propagator in the $k$ sector of a finite system with PBC, provided $L\geq 2r-1$ (i.e., taking a finite system with PBC one selects in a given momentum block only operators with support on $\leq r$ sites). Therefore one can consider our infinite system size $M(k)$ as a specific operator-support selective way of truncating an appropriate unitary matrix. Let us mention at this point that within a random matrix theory one has studied truncations of Haar random unitary matrices, and they result in a spectrum inside a circle whose radius depends on the fraction of retained vectors karol00 .
When numerically dealing with $M(k)$ we never really write out the full matrix (its size $N$ gets too large for large $r$, e.g., $N\approx 5\cdot 10^{7}$ for $r=13$), but rather generate it’s action on the fly by evaluating Eq.(9). This is done by applying each gate of the quantum circuit representation of $U$, see Appendix B.
Due to translational symmetry one also has that $M(-k)$ has the same spectrum as $M(k)$ and therefore one can limit oneself to $k\in[0,\pi)$, rather than $k\in[0,2\pi)$ (the spectrum at $2\pi-k$ is the same as at $k$). In addition to that, due to a spatial reflection symmetry (reversing the order of operators, i.e., in a finite system corresponding to sites mapping $j\to L-j$), which the KI Hamiltonian (I.3) has, the $k=0$ sector splits into two independent parity subsectors corresponding to eigenvalues ±$\pm 1$ of the reflection operator. Namely, $N$ elements of the local basis ${\cal S}_{r}$ can be split into two sets according to whether the reflected operator $Ra$ is the same as $a$, where the reflection of the local $a$ is defined by
@@
with $p$ being the index of the rightmost operator that is not equal to identity, i.e., $a_{p}\in\mathcal{p}$ but $a_{p+1,\ldots,r}=\mathbbm{1}$. That is ${\cal S}_{r}={\cal S}_{r}^{\rm pairs}\cup{\cal S}_{r}^{\rm single}$, where ${\cal S}_{r}^{\rm single}=\{a|(a\in{\cal S}_{r})\,\land\,(Ra=a)\}$, and ${\cal S}_{r}^{\rm pairs}=\{a|(a\in{\cal S}_{r})\,\land\,(Ra\neq a)\}$. Instead of using the basis given by the elements of ${\cal S}_{r}$, one takes an odd basis $\{(a-Ra)/\sqrt{2}|a\in{\cal S}_{r}^{\rm pairs}\}$, and an even basis $\{a|a\in{\cal S}_{r}^{\rm single}\}\cup\{(a+Ra)/\sqrt{2}|a\in{\cal S}_{r}^{\rm pairs}\}$ (where each pair is of course accounted for only once and not twice). The corresponding even and odd sector matrices $M(k=0^{+})$ and $M(k=0^{-})$ are then constructed as in Eqs.(9) and (10) with the only difference being that $a^{\prime}$ and $a$ run over elements of the even or the odd basis, instead of over the full ${\cal S}_{r}$.
As we shall see, studying the momentum dependent RP resonances will give us much more information than just limiting ourselves to $k=0^{+}$ (a sector to which e.g. total magnetization belongs). In particular, different dynamical regimes, like chaotic, mixing, and integrable, will be easier to identify. Note also that some important frequently studied operators are not from the $k=0^{+}$ sector, an example being the magnetization current (i.e., charge current in fermionic language) operator $J$ (the KI model of course does not conserve the magnetization)